Skip to end of metadata
Go to start of metadata

September 17, 2012

Afternoon - Early Evening

Chris and Jesse on site.

Began to dress the Main Tower this afternoon.  Set the Rohn clamps at 20,10,5,4,3,2,1 and 0.5m.  Hung instruments on upper three heights.

20.5m:  0741, B4 (this cable was marked with flagging and a BUNCH of knots.

10.00m:  1119, 1393 (this cable was marked with just flagging tied to coiled cable.

5.01m:  0538, B5 (no cables hung on this level).

4.05m:  

3.03m:

2.0m:

1.04m:

0.50m:

The ATV was unloaded and locked underneath the Base Trailer.  The ramps were also locked to trailer.

The fence stairs have been secured and stable.  A LOT better than first week of set-up.